# Boundary controllability of the Korteweg-de Vries equation: The Neumann case

### v2

@@ 3.2 Construction of the trajectories
The first result of this section ensures that we can construct solutions for the system (3.32) which starts close to 00 (left-hand side of the spatial domain) and achieves some non-null equilibrium in a certain time $T/3$. The result is shown by a fixed-point argument.
@@
where $C>0$ is the positive constant given in Lemma 2.3, [7, Proposition 3.6] guarantees the existence of a bounded linear operator
@@
We will denote, for simplicity, the operator $\Lambda_{1+\varepsilon}$ (given in Proposition 2.4 and Remark 2.9 with $a=1+\varepsilon$) by $\Lambda_{\varepsilon}$. Observe that, if $y$ is solution for (3.32) for some control $h$, then $y$ is a solution of
@@
From Lemma 2.3 and Young inequality we ensure that
@@
and $\delta_{1}$ small enough such that
[…]
@@
The second result of this section ensures the construction of solutions for the system (3.32) on $[2T/3,T]$ starting in one non-null equilibrium and ending near 00.
@@
$C\left(\tau^{\frac{1}{2}}+\tau^{\frac{1}{3}}\right)\|\varepsilon\|_{\mathcal{Z}_{T}}<\delta_{2}\text{ \ \ and \ \ }C\left(\tau^{\frac{1}{2}}+\tau^{\frac{1}{3}}\right)\|\varepsilon\|_{\mathcal{Z}_{T}}^{2}<\delta_{2},$ (3.42)
@@
$=y_{T}-\Lambda_{\varepsilon}(0,0,-zz_{x}+\varepsilon z_{x})(\cdot,\tau)+\Lambda_{\varepsilon}(0,0,-zz_{x}+\varepsilon z_{x})(\cdot,\tau)$
$=y_{T}.$
@@
$y(x,t)=z(x,t-2T/3).$
Now ow, our focus is to show that $\Gamma$ has a fixed point in a suitable metric space. To do that, consider the set $B$ given by
@@
and $\delta_{2}$ small enough such that
[…]
@@
Therefore, $\Gamma:B\rightarrow B$ is a contraction so that, by Banach’s fixed point theorem, $\Gamma$ has a fixed point $z\in B$ which concludes our proof.
@@ 3.3 Controllability on $\mathcal{R}_{c}$
Proof 3.6.
(Proof of Theorem 1.4.) Let $\delta_{1}$ and $\delta_{2}$ be positive real numbers given in Propositions 3.2 and 3.4, respectively. Define $\delta:=\min\{\delta_{1},\delta_{2}\}$ and consider $d\in(0,\delta)$ and $y_{0},y_{T}\in L^{2}(0,L)$ such that
@@
$y=\left\{\begin{array}[]{ll}y_{1},&\text{ in }[0,T/3],\\ d,&\text{ in }[T/3,2T/3],\\ y_{2},&\text{ in }[2T/3,T],\end{array}\right.$ (3.52)
@@ Acknowledgment
Capistrano-Filho was supported by CAPES grant number 88881.311964/2018-01, COFECUB/CAPES grant number 8887.879175/2023-00, CNPq grants numbers 307808/2021-1 and 401003/2022-1. Da Silva acknowledges support from CNPq. This work is part of the Ph.D. thesis of da Silva at the Department of Mathematics of the Universidade Federal de Pernambuco.

### v3

@@ 3.2 Construction of the trajectories
The first result of this section ensures that we can construct solutions for the system (3.32) which starts close to 0 (left-hand side of the spatial domain) and achieves some non-null equilibrium in a certain time $T/3$. The result is shown by a fixed-point argument.
@@
where $C>0$ is the positive constant given in Lemma 2.3, [7, Proposition 3.6] guarantees the existence of a bounded linear control operator
@@
We will denote, for simplicity, the solution operator $\Lambda_{1+\varepsilon}$ (given in Proposition 2.4 and Remark 2.9 with $a=1+\varepsilon$) by $\Lambda_{\varepsilon}$. Observe that, if $y$ is solution for (3.32) for some control $h$, then $y$ is a solution of
@@
From Lemma 2.3 and Young’s inequality, we ensure that
@@
and $\delta_{1}$ is small enough such that
$\left(C_{2}^{2}\|\Psi^{\varepsilon}\|+C_{2}\right)\overline{C}r<\frac{1}{2},\ \ \ 2\left(C_{2}^{2}\|\Psi^{\varepsilon}\|+C_{2}\right)r<\frac{1}{2},\ \ \ \left(C_{2}^{2}\|\Psi^{\varepsilon}\|+C_{2}\right)\delta_{1}<\frac{1}{2},$ (3.41)
@@
The second result of this section ensures the construction of solutions for the system (3.32) on $[2T/3,T]$ starting in one non-null equilibrium and ending near 0.
@@
[…]
@@
$=y_{T}-\Lambda_{\varepsilon}(0,0,-zz_{x}+\varepsilon z_{x})(\cdot,\tau)+\Lambda_{\varepsilon}(0,0,-zz_{x}+\varepsilon z_{x})(\cdot,\tau)=y_{T}.$
@@
y(x,t)=z(x,t2T/3).y(x,t)=z(x,t-2T/3).
Now, our focus is to show that $\Gamma$ has a fixed point in a suitable metric space. To do that, consider the set $B$ given by
@@
and $\delta_{2}$ is small enough such that
$\left(C_{2}^{2}\|\Psi^{\varepsilon}\|+C_{2}\right)\overline{C}r<\frac{1}{2},\ \ \ 2\left(C_{2}^{2}\|\Psi^{\varepsilon}\|+C_{2}\right)r<\frac{1}{2},\ \ \ \left(C_{2}^{2}\|\Psi^{\varepsilon}\|+C_{2}\right)\delta_{1}<\frac{1}{2},$ (3.48)
@@
Therefore, $\Gamma:B\rightarrow B$ is a contraction so that, by Banach’s fixed point theorem, $\Gamma$ has a fixed point $z\in B$, which concludes our proof.
Remark 3.7.
It is important to note that Propositions 3.3 and 3.5 guarantee that $\|\Psi^{\varepsilon}\|$ can be chosen uniformly bounded for sufficiently small $\varepsilon$. This uniform bound is crucial to ensure the existence of a fixed point in both theorems.
@@ 3.3 Controllability on $\mathcal{R}_{c}$
Proof 3.8.
(Proof of Theorem 1.4.) Let $\delta_{1}$ and $\delta_{2}$ be positive real numbers given in Propositions 3.3 and 3.5, respectively. Define $\delta:=\min\{\delta_{1},\delta_{2}\}$ and consider $d\in(0,\delta)$ and $y_{0},y_{T}\in L^{2}(0,L)$ such that
@@
$y=\left\{\begin{array}[]{ll}y_{1},&\text{ in }[0,T/3],\\ d,&\text{ in }[T/3,2T/3],\\ y_{2},&\text{ in }[2T/3,T],\end{array}\right$.
@@
4 Conclusion and open issues
In this work, we established the exact boundary controllability of the nonlinear Korteweg–de Vries equation under Neumann boundary conditions in the critical length case. By combining the return method of Coron [14] with a suitable fixed-point argument, we extended previous results in the literature that were restricted to the subcritical regime. Our analysis shows that, for small perturbations of the steady state $u\equiv c$, the system is exactly controllable in $L^{2}(0,L)$ when $L\in\mathcal{R}_{c}$. This completes the controllability theory for the Neumann case and opens perspectives for future research on control and several important conclusions, which we will mention now.
4.1 Control cost and small-time behavior
Observe that Theorems 1.4 and 1.5, and consequently Theorem 1.6, guarantee controllability in small time. This is in contrast with the setting considered in [7, 24], where the control cost cannot be directly characterized. In our approach, small-time controllability follows from Propositions 3.3 and 3.5, specifically from the assumptions on $\tau$ given in (3.39) and (3.42).
This naturally leads to the question of characterizing the constant in the observability inequality associated with the control problem (1.9). Moreover, we expect the control $h$ to blow up in the $L^{2}$–norm as $T\to 0^{+}$, since the control is bounded by the constant
$\overline{C}:=\frac{3C}{2}\left(T^{\frac{1}{2}}+T^{\frac{1}{3}}\right).$
Both questions remain open and provide interesting directions for future research on the control problem.
4.2 Neumann versus Dirichlet/mixed controls
In this work, we investigated the Korteweg–de Vries (KdV) equation on a bounded domain under Neumann boundary conditions. In the classical Hilbert Uniqueness Method (HUM), the main difficulty lies in the fact that the boundary conditions for the adjoint system differ from those of the original system. For this reason, the return method cannot be applied directly, unlike approaches based on series expansions as in the works of Crépeau, Cerpa, Coron, and others [11, 10]. To overcome this issue, we adapted Glass’s idea [16] and employed a fixed-point argument applied to a specifically constructed trajectory, which is, in general, not unique.
Several open problems remain concerning the KdV equation with Neumann boundary conditions. For instance, it is still unclear how to use power series expansions to obtain similar results as in [11] or how to define a manifold (a subset of $L^{2}(0,L)$) associated with the set $\mathcal{R}_{c}$ to characterize all admissible trajectories in the critical length regime, see for instance the references [8, 10]. In this sense, the present work not only extends the existing literature but also highlights important directions for future research.
@@ Acknowledgment
The authors are grateful to the referee for the careful reading of this paper and their valuable suggestions and comments. Capistrano-Filho was supported by CAPES grant number 88881.311964/2018-01, COFECUB/CAPES grant number 8887.879175/2023-00, CNPq grants numbers 307808/2021-1 and 401003/2022-1. Da Silva acknowledges support from CNPq. This work is part of the Ph.D. thesis of da Silva at the Department of Mathematics of the Universidade Federal de Pernambuco.